OpenStudy (anonymous):

what number completes the pattern 1,3,4,?,11,18,29

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

do u guys get it?

OpenStudy (aripotta):

7

OpenStudy (anonymous):

This is called a Fibonacci Sequence. Look at the first 3 numbers, 1,3,4. What you notice is if you add the first two numbers, you get the third number. 1+3=4. now Look at the last 3 numbers, once again, 11+18=29. so to find the missing number take the 18, and subtract 11 from it.

OpenStudy (anonymous):

In a sequence like this, any 3 given digits, the first two added together will equal the third. get it?

OpenStudy (anonymous):

so i add 3+4=7 which is my answer?

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok thank you very much